# On uniqueness of radial potentials for given Dirichlet spectra with distinct angular momenta

### v1

@@ 8 Uniqueness result in the case $\ell=0$ and $\ell=2$
In particular, this implies that the spectral map associated with the pair $(\ell_{1},\ell_{2})=(0,2)$ is injective.
9 Conclusion
The techniques developed in this paper rely exclusively on the Kneser–Sommerfeld formula. Even with computer algebra, the analysis of the resulting differential equations is already cumbersome in the cases $(\ell_{1},\ell_{2})=(1,2)$ and $(0,3)$, and larger values currently seem out of reach. This suggests that a more conceptual approach is needed.
Appendix A A Bessel basis of $L^{2}(0,1)$
In [23, Proposition 6.1], Rundell and Sacks established the completeness in $L^{2}(0,1)$ of the family
$\Bigl(1,\ \bigl(\Phi_{0}(n\pi\cdot)\bigr)_{n\geq 1},\ \bigl(\Phi_{2}((n+1)\pi\cdot)\bigr)_{n\geq 1}\Bigr),$ (A.1)
where, for any integer $\ell$,
$\Phi_{\ell}(x):=\tfrac{\pi}{2}\,x\,J_{\ell+\tfrac{1}{2}}(x)^{2}.$
The aim of this appendix is to strengthen their result by proving that the very same family does not merely provide completeness, but in fact constitutes a basis of $L^{2}(0,1)$.
The next lemma will play a key role in the proof of Theorem A.2.
Lemma A.1.
Let $f\in L^{2}(0,1)$. Suppose that both $f$ and $g:=T_{2}^{*}f$ are even with respect to $\tfrac{1}{2}$. Moreover, suppose that $f$ is orthogonal to the constants and to $\cos(2\pi x)$. Then $f=0$.
Proof.
Using Lemma 4.2 successively for $\ell=1$ and $\ell=2$, and recalling that $T_{2}^{*}=-S_{2}^{*}S_{1}^{*}$, we obtain:
$g^{(5)}+\frac{12}{x}\,g^{(4)}+\frac{24}{x^{2}}\,g^{(3)}-\frac{24}{x^{3}}\,g^{\prime\prime}\;=\;-f^{(5)}.$ (A.2)
Projecting this equation onto the subspace of even functions and setting $y=g^{\prime\prime}$, we obtain the following ordinary differential equation:
$y^{\prime\prime}+2\,\frac{1-2x}{x(1-x)}\,y^{\prime}-2\,\frac{3x^{2}-3x+1}{x^{2}(1-x)^{2}}\,y=0.$ (A.3)
The roots of the indicial equation associated with (A.3) are $r=1$ and r=2r=-2. Actually, the general solution of (A.3) on $(0,1)$ can be explicitly written in terms of Ferrers functions (see [18], 14.2) as
$y(x)=g^{\prime\prime}(x)=\frac{C_{1}\,\mathbf{P}^{\,3}_{\nu}(2x-1)+C_{2}\,\mathbf{Q}^{\,3}_{\nu}(2x-1)}{\sqrt{x(1-x)}},\qquad\nu=\frac{\sqrt{33}-1}{2}.$ (A.4)
To enforce the symmetry y(1x)=y(x)y(1-x)=y(x), the coefficients must satisfy (see [18], Section 14.3)
$\frac{C_{2}}{C_{1}}\;=\;\frac{2}{\pi}\,\cot\!\Big(\frac{\pi\nu}{2}\Big).$
An even solution is therefore of the form $y(x)=K\,y_{0}(x)$ for some constant $K\in\mathbb{C}$, where
$y_{0}(x)=\frac{\mathbf{P}^{\,3}_{\nu}(2x-1)\;+\;\tfrac{2}{\pi}\,\cot\!\left(\tfrac{\pi\nu}{2}\right)\,\mathbf{Q}^{\,3}_{\nu}(2x-1)}{\sqrt{x(1-x)}}.$ (A.5)
Using (A.2), we obtain
$y^{(3)}(x)+\frac{12}{x}\,y^{\prime\prime}(x)+\frac{24}{x^{2}}\,y^{\prime}(x)-\frac{24}{x^{3}}\,y(x)\;=\;-\,f^{(5)}(x).$ (A.6)
To simplify notation, set
$z_{0}(x):=y_{0}^{(3)}(x)+\frac{12}{x}\,y_{0}^{\prime\prime}(x)+\frac{24}{x^{2}}\,y_{0}^{\prime}(x)-\frac{24}{x^{3}}\,y_{0}(x).$ (A.7)
Consequently, since $f$ is even with respect to $x=\tfrac{1}{2}$, we may write
$f(x)\;=\;K\,f_{0}(x)\;+\;P(x),$ (A.8)
where
$f_{0}(x)\;=\;-\,\frac{1}{24}\int_{1/2}^{x}(x-t)^{4}\,z_{0}(t)\,dt,$ (A.9)
and
$P(x)\;=\;a\;+\;b\,(x-\tfrac{1}{2})^{2}\;+\;c\,(x-\tfrac{1}{2})^{4}.$
Here $a$, $b$, and $c$ are constants determined by the even Taylor expansion of $f$ at $x=\tfrac{1}{2}$:
$a\;=\;f\!\left(\tfrac{1}{2}\right),\qquad b\;=\;\tfrac{1}{2}\,f^{\prime\prime}\!\left(\tfrac{1}{2}\right),\qquad c\;=\;\tfrac{1}{24}\,f^{(4)}\!\left(\tfrac{1}{2}\right).$ (A.10)
Now, using the definition of $T_{2}^{*}$,
$g(x)=T_{2}^{*}f(x)\;=\;-f(x)-\frac{12}{x^{2}}\int_{0}^{x}tf(t)\,dt\;+\;\frac{24}{x^{4}}\int_{0}^{x}t^{3}f(t)\,dt,$ (A.11)
we readily obtain
$(g+f)^{\prime}(x)$ $=\frac{12}{x}\,f(x)\;+\;\frac{24}{x^{3}}\int_{0}^{x}t\,f(t)\,dt\;-\;\frac{96}{x^{5}}\int_{0}^{x}t^{3}f(t)\,dt,$ (A.12)
$(g+f)^{(3)}(x)$ $=\frac{12}{x}\,f^{\prime\prime}(x)\;-\;\frac{96}{x^{2}}\,f^{\prime}(x)\;+\;\frac{576}{x^{3}}\,f(x)$
$\quad+\;\frac{288}{x^{5}}\int_{0}^{x}t\,f(t)\,dt\;-\;\frac{2880}{x^{7}}\int_{0}^{x}t^{3}f(t)\,dt.$ (A.13)
Consequently, since $f$ and $g$ are even, we obtain at $x=\frac{1}{2}$,
$0\;=\;(g+f)^{\prime}\!\left(\tfrac{1}{2}\right)$ $=24\,a\;+\;192\int_{0}^{1/2}t\,f(t)\,dt\;-\;3072\int_{0}^{1/2}t^{3}f(t)\,dt,$ (A.14)
$0\;=\;(g+f)^{(3)}\!\left(\tfrac{1}{2}\right)$ $=4608\,a\;+\;48\,b\;+\;9216\int_{0}^{1/2}t\,f(t)\,dt\;-\;368640\int_{0}^{1/2}t^{3}f(t)\,dt.$ (A.15)
Substituting $f=Kf_{0}+P$, a straightforward calculation yields:
$b\;+\;\tfrac{1}{7}\,c$ $=K\!\left(-\,960\int_{0}^{1/2}t\,f_{0}(t)\,dt\;+\;15360\int_{0}^{1/2}t^{3}f_{0}(t)\,dt\right),$ (A.16)
$c$ $=K\!\left(26880\int_{0}^{1/2}t\,f_{0}(t)\,dt\;-\;1\,075\,200\int_{0}^{1/2}t^{3}f_{0}(t)\,dt\right).$ (A.17)
From (A.16) and (A.17), we deduce that
$b\;=\;960\,K\!\left(-\,5\int_{0}^{1/2}t\,f_{0}(t)\,dt\;+\;176\int_{0}^{1/2}t^{3}f_{0}(t)\,dt\right).$ (A.18)
Moreover, since $f$ is orthogonal to $\cos(2\pi x)$, we obtain
$b\;+\;\frac{\pi^{2}-6}{2\pi^{2}}\,c\;=\;-\,2\pi^{2}K\int_{0}^{1}f_{0}(x)\,\cos(2\pi x)\,dx.$ (A.19)
A numerical computation with Mathematica yields

### v3

@@ 8 Uniqueness result in the case $(\ell_{1},\ell_{2})=(0,2)$
In particular, this implies that the differential of the spectral map associated with the pair $(\ell_{1},\ell_{2})=(0,2)$ is injective.
9 Uniqueness result in the case $(\ell_{1},\ell_{2})=(1,2)$
In this section, we assume that $\zeta$ satisfies the hypotheses of Proposition 5.1 for both $\ell=1$ and $\ell=2\,$. Our goal is to prove that $\zeta$ vanishes identically on (0,1)(0,1)\,.
We first consider the case $\ell=1\,$. By Theorem 6.6, the distribution $A_{1}(D)\bigl[S_{1}[\zeta]\bigr]$ is odd, where $A_{1}(t)=1-\frac{t}{2}$ and $D=\frac{d}{dx}\,$. Thus, if we denote $f:=S_{1}[\zeta]\,$, this condition is equivalent to requiring that $f^{\prime}-2f$ be odd (in the sense of distributions). Decomposing $f$ into its even and odd parts,
$f=f_{e}+f_{o}\,$, (9.1)
where $f_{e}$ is even and $f_{o}$ is odd with respect to $x=\tfrac{1}{2}$ , we immediately get
$f_{e}=\frac{1}{2}f_{o}^{\prime}\,$, (9.2)
and therefore
$f=\frac{1}{2}f_{o}^{\prime}+f_{o}\,.$ (9.3)
We now exploit the case $\ell=2\,$. Since $\zeta$ satisfies the assumptions of Proposition 5.1 with $\ell=2$, Theorem 6.6 yields that $A_{2}(D)\bigl[T_{2}[\zeta]\bigr]$ is odd in the sense of distributions, where
$A_{2}(t)=\tfrac{1}{4}t^{2}-\tfrac{3}{2}t+3\,.$
We define
$g:=S_{2}[f]=f(x)-8x^{3}\int_{x}^{1}\frac{f(t)}{t^{4}}\,dt\,$,
so that $g=-T_{2}[\zeta]\,$. A straightforward computation yields
$4A_{2}(D)[g](x)=f^{\prime\prime}(x)+\Bigl(\frac{8}{x}-6\Bigr)f^{\prime}(x)+\Bigl(\frac{16}{x^{2}}-\frac{48}{x}+12\Bigr)f(x)-48x(1-x)(1-2x)\int_{x}^{1}\frac{f(t)}{t^{4}}\,dt\,.$ (9.4)
Since $4A_{2}(D)[g]$ is odd with respect to $x=\tfrac{1}{2}$ , evaluating (9.4) at $x=\tfrac{1}{2}$ yields
$f^{\prime\prime}\!\left(\tfrac{1}{2}\right)+10\,f^{\prime}\!\left(\tfrac{1}{2}\right)-20\,f\!\left(\tfrac{1}{2}\right)=0\,.$
Using the decomposition $f=\tfrac{1}{2}f_{o}^{\prime}+f_{o}$, where $f_{o}$ is odd, we obtain
$f_{o}^{(3)}\!\left(\tfrac{1}{2}\right)=0\,.$ (9.5)
Similarly, a straightforward computation gives
[…]
$\quad+\Bigl(\frac{96}{x}-\frac{48}{x^{2}}\Bigr)f^{\prime}(x)+\Bigl(\frac{192}{x^{2}}-\frac{96}{x^{3}}\Bigr)f(x)$
[…]
Since $D^{2}\bigl(4A_{2}(D)[g]\bigr)$ is odd with respect to $x=\tfrac{1}{2}$, evaluating at $x=\tfrac{1}{2}$ gives
$f_{o}^{(5)}\!\left(\tfrac{1}{2}\right)=64\,f_{o}^{(3)}\!\left(\tfrac{1}{2}\right)=0\,.$ (9.7)
Finally, a similar computation yields
[…]
$\quad+\Bigl(\frac{96}{x}+\frac{48}{x^{2}}+\frac{16}{x^{3}}\Bigr)f^{(3)}(x)\,.$
Replacing, $f=\frac{1}{2}f_{o}^{\prime}+f_{o}$, we get
$G(x)$ $=\frac{1}{2}\,f_{o}^{(7)}(x)+\Bigl(\frac{4}{x}-2\Bigr)f_{o}^{(6)}(x)-\Bigl(\frac{16}{x}+\frac{8}{x^{2}}\Bigr)f_{o}^{(5)}(x)$ (9.9)
[…]
We now use the symmetry condition G(x)+G(1x)=0G(x)+G(1-x)=0. This yields the following differential equation.
$0={}$ $f_{o}^{(7)}+4\left(\frac{1}{x}-\frac{1}{1-x}\right)f_{o}^{(6)}$ (9.10)
$-\left[16\left(\frac{1}{x}+\frac{1}{1-x}\right)+8\left(\frac{1}{x^{2}}+\frac{1}{(1-x)^{2}}\right)\right]f_{o}^{(5)}$
$+8\left[\left(\frac{1}{x^{2}}-\frac{1}{(1-x)^{2}}\right)+\left(\frac{1}{x^{3}}-\frac{1}{(1-x)^{3}}\right)\right]f_{o}^{(4)}$
$+\left[96\left(\frac{1}{x}+\frac{1}{1-x}\right)+48\left(\frac{1}{x^{2}}+\frac{1}{(1-x)^{2}}\right)+16\left(\frac{1}{x^{3}}+\frac{1}{(1-x)^{3}}\right)\right]f_{o}^{(3)}\,.$
Finally, setting
$y(x)=f_{o}^{(3)}(x)\,$,
we obtain a fourth–order differential equation for $y$ . The function $y$ is even with respect to $x=\tfrac{1}{2}$, and from the previous identities we have
$y^{(k)}\!\left(\tfrac{1}{2}\right)=0\,,\qquad k=0,1,2,3\,.$
By the Cauchy–Lipschitz theorem, it follows that $y\equiv 0$. Hence $f_{o}$ is an odd polynomial of degree at most two, and therefore necessarily of the form
$f_{o}(x)=a\,(x-\tfrac{1}{2})\,.$ (9.11)
Using again $f=\tfrac{1}{2}f_{o}^{\prime}+f_{o}\,$, we obtain
$f(x)=\frac{a}{2}+a\,(x-\tfrac{1}{2})=a\,x\,.$ (9.12)
Since $f=S_{1}[\zeta]$, applying the left inverse $A_{1}$ of $S_{1}$ given in Lemma 4.1(iii) yields
$\zeta(x)=A_{1}[f](x)=f(x)-\frac{4}{x^{3}}\int_{0}^{x}t^{2}\,f(t)\,dt=0\,.$ (9.13)
We have thus proved the following theorem.
Theorem 9.1.
Assume that $\zeta\in L^{2}(0,1)$ satisfies the hypotheses of Proposition 5.1 for $\ell=1$ and $\ell=2$. Then $\zeta=0$ almost everywhere on $(0,1)$ .
In particular, the differential of the spectral map associated with the pair $(\ell_{1},\ell_{2})=(1,2)$ is injective.
10 Uniqueness result in the case $(\ell_{1},\ell_{2})=(0,3)$
Throughout this section, we assume that $\zeta$ satisfies the hypotheses of Proposition 5.1 for both $\ell=0$ and $\ell=3$ , and we aim to show that $\zeta$ vanishes almost everywhere on $(0,1)$ . As before, the condition for $\ell=0$ implies that $\zeta$ is odd with respect to the midpoint $x=\tfrac{1}{2}$ . For $\ell=3$ , Theorem 6.6 implies that the function $A_{3}(D)\bigl[T_{3}[\zeta]\bigr]$ is odd in the sense of distributions.
In order to prove that $\zeta$ vanishes almost everywhere on $(0,1)$, we follow the same strategy as in the previous section. The only step where we rely on computer assistance is Step 3, where we provide a computer-assisted proof.
  • Step 1. We show that
    $\Bigl(\forall\,k\in\{0,1,2\},\ D^{2k}A_{3}(D)\bigl[T_{3}[\zeta]\bigr]\!\left(\tfrac{1}{2}\right)=0\Bigr)\quad\Longrightarrow\quad(S)\,$,
    where $(S)$ denotes a triangular linear system involving the even-order derivatives of $y=\zeta^{\prime}$ at $x=\tfrac{1}{2}$ , uniquely determined by the single value $y(\tfrac{1}{2})$ .
  • Step 2. We show that $y$ satisfies a linear differential equation of order $8$ on the interval $(0,1)$. The associated indicial equation at $x=0$ has the characteristic roots
    $\rho=-2,\;1,\;3,\;5,\;6,\;7,\;-2\pm 2i\sqrt{11}\,.$
  • Step 3. A numerical investigation shows that the unique solution of this differential equation satisfying $y(\tfrac{1}{2})=1$ and whose initial data at $x=\tfrac{1}{2}$ are prescribed by the system $(S)$ exhibits a pronounced blow-up as $x\to 0^{+}$ and $x\to 1^{-}$. The numerical solution $y(x)$ displays oscillatory growth compatible with the Frobenius exponents $\alpha=-2\pm 2i\sqrt{11}$ . As a consequence, the corresponding function $\zeta$ fails to be square-integrable on $(0,1)$ .
  • Step 4. We infer that all derivatives of $y$ at $x=\tfrac{1}{2}$ must vanish. It then follows that $y\equiv 0$ on $(0,1)$, and consequently $\zeta=0$ almost everywhere on $(0,1)$, since $\zeta$ is odd with respect to the midpoint $x=\tfrac{1}{2}$ .
Step 1.a: for all odd $k\in\{0,\dots 7\}$, i.e., $k\in\{1,3,5,7\}$  , $y^{(k)}\!\left(\tfrac{1}{2}\right)=0$ . This is an immediate consequence of the fact that $y^{(k)}$ is odd whenever $k$ is odd.
Step 1.b. We derive a triangular linear system $(S)$ for the even-order jets of $y$ at $x=\tfrac{1}{2}$ . We follow the strategy used in cases $(0,1)$ and $(0,2)$ but as we will see, due to the increasing degree of the polynomial $A_{\ell}$ , we need to work a little more.
Step 1.b.i: using $A_{3}(D)\bigl[T_{3}[\zeta]\bigr]\!\left(\tfrac{1}{2}\right)=0$. Recall that (see (3.9) in [27]), for all $x\in(0,1)$ ,
$h(x)=T_{3}[\zeta](x)=\zeta(x)-24x\int_{x}^{1}\frac{\zeta(t)}{t^{2}}\,dt+120x^{3}\int_{x}^{1}\frac{\zeta(t)}{t^{4}}\,dt-120x^{5}\int_{x}^{1}\frac{\zeta(t)}{t^{6}}\,dt\,.$
Applying the differential operator $8A_{3}(D)=-D^{3}+12D^{2}-60D+120$ to $h$, we obtain for all $x\in(0,1)$ ,
$8A_{3}(D)[h](x)=$ $-\zeta^{\prime\prime\prime}(x)+\left(12-\frac{24}{x}\right)\zeta^{\prime\prime}(x)+\left(-60+\frac{288}{x}-\frac{216}{x^{2}}\right)\zeta^{\prime}(x)$
$+\left(120-\frac{1440}{x}+\frac{2880}{x^{2}}-\frac{1200}{x^{3}}\right)\zeta(x)+1440(-2x+1)\int_{x}^{1}\frac{\zeta(t)}{t^{2}}\,dt$
$+720(20x^{3}-30x^{2}+12x-1)\int_{x}^{1}\frac{\zeta(t)}{t^{4}}\,dt$
$+7200(-2x^{5}+5x^{4}-4x^{3}+x^{2})\int_{x}^{1}\frac{\zeta(t)}{t^{6}}\,dt\,.$
Evaluating this expression at $x=\tfrac{1}{2}$ and using that $\zeta$ is odd and $y=\zeta^{\prime}$, we obtain
$8A_{3}(D)\bigl[h\bigr]\!\left(\tfrac{1}{2}\right)=-\zeta^{\prime\prime\prime}\!\left(\tfrac{1}{2}\right)-36\zeta^{\prime\prime}\!\left(\tfrac{1}{2}\right)-348\zeta^{\prime}\!\left(\tfrac{1}{2}\right)-840\zeta\!\left(\tfrac{1}{2}\right)=-y^{\prime\prime}\!\left(\tfrac{1}{2}\right)-348\,y\!\left(\tfrac{1}{2}\right)\,.$
Since $8A_{3}(D)[h]$ is odd, we therefore conclude that
$y^{\prime\prime}\!\left(\tfrac{1}{2}\right)-348\,y\!\left(\tfrac{1}{2}\right)=0\,.$
Step 1.b.ii: using $D^{2}A_{3}(D)\bigl[T_{3}[\zeta]\bigr]\!\left(\tfrac{1}{2}\right)=0$. After some calculations we obtain :
$8D^{2}A_{3}(D)\bigl[h\bigr]\!\left(\tfrac{1}{2}\right)=-\zeta^{(5)}\!\left(\tfrac{1}{2}\right)-36\zeta^{(4)}\!\left(\tfrac{1}{2}\right)-156\zeta^{\prime\prime\prime}\!\left(\tfrac{1}{2}\right)+3384\zeta^{\prime\prime}\!\left(\tfrac{1}{2}\right)+18432\zeta^{\prime}\!\left(\tfrac{1}{2}\right)\,.$
Since $8D^{2}A_{3}(D)[h]$ is odd, we therefore conclude that
$y^{(4)}\!\left(\tfrac{1}{2}\right)-156y^{\prime\prime}\!\left(\tfrac{1}{2}\right)+18432\,y\!\left(\tfrac{1}{2}\right)=0\,.$
Step 1.b.iii: using $D^{4}A_{3}(D)\bigl[T_{3}[\zeta]\bigr]\!\left(\tfrac{1}{2}\right)=0$. After some calculations we obtain :
$8D^{4}A_{3}(D)\bigl[h\bigr]\!\left(\tfrac{1}{2}\right)=-\zeta^{(7)}\!\left(\tfrac{1}{2}\right)-36\zeta^{(6)}\!\left(\tfrac{1}{2}\right)+36\zeta^{(5)}\!\left(\tfrac{1}{2}\right)+6072\zeta^{(4)}\!\left(\tfrac{1}{2}\right)-18432\zeta^{\prime\prime\prime}\!\left(\tfrac{1}{2}\right)-294912\zeta^{\prime\prime}\!\left(\tfrac{1}{2}\right)+589824\zeta^{\prime}\!\left(\tfrac{1}{2}\right)\,.$
Since $8D^{4}A_{3}(D)[h]$ is odd, we therefore conclude that
$y^{(6)}\!\left(\tfrac{1}{2}\right)+36y^{(4)}\!\left(\tfrac{1}{2}\right)-18432y^{\prime\prime}\!\left(\tfrac{1}{2}\right)+589824y\!\left(\tfrac{1}{2}\right)=0\,.$
Thus, in contrast with the case $(0,2)$, where for $k\in\{0,1\}$ one has
$D^{2k}A_{2}(D)\bigl[T_{2}[\zeta]\bigr]\!\left(\tfrac{1}{2}\right)=0\quad\Longrightarrow\quad y^{(2k)}\!\left(\tfrac{1}{2}\right)=0\,$,
we obtain here
$\Bigl(\forall\,k\in\{0,1,2\}\,,\ D^{2k}A_{3}(D)\bigl[T_{3}[\zeta]\bigr]\!\left(\tfrac{1}{2}\right)=0\Bigr)\quad\Longrightarrow\quad(S)\,$,
where $(S)$ denotes the following triangular linear system:
$\left\{\begin{aligned} &&&&&&&&&&&y^{\prime\prime}\!\left(\tfrac{1}{2}\right)&{}+{}&348&\,y\!\left(\tfrac{1}{2}\right)&=0,\\ &&&&&&&&y^{(4)}\!\left(\tfrac{1}{2}\right)&{}+{}&156&\,y^{\prime\prime}\!\left(\tfrac{1}{2}\right)&{}-{}&18432&\,y\!\left(\tfrac{1}{2}\right)&=0,\\ &&&&&y^{(6)}\!\left(\tfrac{1}{2}\right)&{}-{}&36&\,y^{(4)}\!\left(\tfrac{1}{2}\right)&{}+{}&18432&\,y^{\prime\prime}\!\left(\tfrac{1}{2}\right)&{}-{}&589824&\,y\!\left(\tfrac{1}{2}\right)&=0\,.\end{aligned}\right$.
This system is uniquely determined by the single value $y(\tfrac{1}{2})$ .
Step 2 : $y$ is a solution of a linear $8^{\text{th}}$-order differential equation. Let us set $f:=S_{1}[\zeta]$ , $g:=S_{2}[f]$ , and $h:=S_{3}[g]$, so that $h=T_{3}[\zeta]=S_{3}S_{2}S_{1}[\zeta]$ . Using Lemma 4.1 (iv) with $\ell=3$, $\ell=2$ and $\ell=1$ , we have
$D^{6}[h]=D^{6}[g]+\frac{12}{x}D^{5}[g]\quad;\quad D^{4}[g]=D^{4}[f]+\frac{8}{x}D^{3}[f]\quad;\quad D^{2}[f]=D^{2}[\zeta]+\frac{4}{x}D[\zeta].$
Then, we can write $D^{6}[h]$ as
$D^{6}[h]=D^{6}[\zeta]+\sum_{k=1}^{5}a_{k}\!\left(\tfrac{1}{x}\right)D^{k}[\zeta]$ (10.1)
where for all $k\in\{1,\dots,5\}\,$, $a_{k}$ is a polynomial function. We now differentiate the expression $A_{3}(D)\bigl[h\bigr]$ six times. Since applying an even number of derivatives preserves parity in the distributional sense, and since the operator $A_{3}(D)$ commutes with differentiation, it follows that the function
$G:=8D^{6}A_{3}(D)\bigl[h\bigr]=8A_{3}(D)\bigl[D^{6}[h]\bigr]$
is odd. Using the expression of $A_{3}(D)$ given in Remark 6.7 and the form of $D^{6}[h]$ given in (10.1), we obtain that the odd function $G$ can be written as
$G(x)=(-D^{3}+12D^{2}-60D+120)\bigl[D^{6}[\zeta]\bigr](x)+\sum_{k=1}^{5}a_{k}\!\left(\tfrac{1}{x}\right)D^{k}[\zeta](x)=-D_{9}[\zeta](x)+\sum_{k=1}^{8}b_{k}\!\left(\tfrac{1}{x}\right)D^{k}[\zeta](x)\,$,
where for all $k\in\{1,\dots,8\}$, $b_{k}$ is a polynomial function. Thus, since $G$ is odd on $(0,1)$, it satisfies G(x)+G(1x)=0G(x)+G(1-x)=0 in the sense of distributions. It follows that $y=\zeta^{\prime}=D[\zeta]$ satisfies the following linear $8^{\text{th}}$-order differential equation on $(0,1)$:
Lemma 10.1.
Assume that $\zeta\in L^{2}(0,1)$ satisfies the hypotheses of Lemma 5.1 for $\ell=0$ and $\ell=3$ , and set $y(x):=\zeta^{\prime}(x)$ . Then $y$ satisfies the following differential equation
$y^{(8)}(x)+12\left(\frac{1}{1-x}-\frac{1}{x}\right)y^{(7)}(x)$
$\left[-60+144\left(\frac{1}{x}+\frac{1}{1-x}\right)-36\left(\frac{1}{x^{2}}+\frac{1}{(1-x)^{2}}\right)\right]y^{(6)}(x)$
$\left[-720\left(\frac{1}{x}-\frac{1}{1-x}\right)+576\left(\frac{1}{x^{2}}-\frac{1}{(1-x)^{2}}\right)+336\left(\frac{1}{x^{3}}-\frac{1}{(1-x)^{3}}\right)\right]y^{(5)}(x)$
$\left[-720\left(\frac{1}{x^{4}}+\frac{1}{(1-x)^{4}}\right)-2880\left(\frac{1}{x^{3}}+\frac{1}{(1-x)^{3}}\right)-3600\left(\frac{1}{x^{2}}+\frac{1}{(1-x)^{2}}\right)+1440\left(\frac{1}{x}+\frac{1}{1-x}\right)\right]y^{(4)}(x)$
$\left[-2880\left(\frac{1}{x^{5}}-\frac{1}{(1-x)^{5}}\right)+7200\left(\frac{1}{x^{3}}-\frac{1}{(1-x)^{3}}\right)+8640\left(\frac{1}{x^{2}}-\frac{1}{(1-x)^{2}}\right)\right]y^{(3)}(x)$
$\left[23040\left(\frac{1}{x^{6}}+\frac{1}{(1-x)^{6}}\right)+34560\left(\frac{1}{x^{5}}+\frac{1}{(1-x)^{5}}\right)+21600\left(\frac{1}{x^{4}}+\frac{1}{(1-x)^{4}}\right)+2880\left(\frac{1}{x^{3}}+\frac{1}{(1-x)^{3}}\right)\right]y^{\prime\prime}(x)$
$\left[60480\left(\frac{1}{x^{7}}-\frac{1}{(1-x)^{7}}\right)+103680\left(\frac{1}{x^{6}}-\frac{1}{(1-x)^{6}}\right)+86400\left(\frac{1}{x^{5}}-\frac{1}{(1-x)^{5}}\right)+34560\left(\frac{1}{x^{4}}-\frac{1}{(1-x)^{4}}\right)\right]y^{\prime}(x)$
$\left[60480\left(\frac{1}{x^{8}}+\frac{1}{(1-x)^{8}}\right)+103680\left(\frac{1}{x^{7}}+\frac{1}{(1-x)^{7}}\right)+86400\left(\frac{1}{x^{6}}+\frac{1}{(1-x)^{6}}\right)+34560\left(\frac{1}{x^{5}}+\frac{1}{(1-x)^{5}}\right)\right]y(x)\;=\;0\,$,
in the sense of distributions on $(0,1)$ .
The indicial equation at $x=0$ associated with the differential equation of Lemma 10.1 takes the form
$(\rho-7)(\rho-6)(\rho-5)(\rho-3)(\rho-1)(\rho+2)\left(\rho^{2}+4\rho+48\right)=0\,.$
Consequently, the corresponding indicial roots are
$\rho\;=\;-2,\;1,\;3,\;5,\;6,\;7,\;-2\pm 2i\sqrt{11}\,.$
Step 3: Numerical study of the solution. We numerically integrate the eighth–order differential equation given in Lemma 10.1 with initial conditions prescribed at $x=\tfrac{1}{2}$ , namely $y(\tfrac{1}{2})=1$ together with the even derivatives $y^{(2k)}(\tfrac{1}{2})$ computed in Step 1, the odd derivatives being zero by symmetry. The resulting solution $y$, which is even with respect to $x=\tfrac{1}{2}$ , blows up at both endpoints $x=0$ and $x=1$ .
[…]
Guided by these indicial roots, we extract the boundary asymptotics at $x=0$ by fitting $x^{2}y(x)$ on shrinking windows near the boundary. We obtain numerically
$y(x)=\frac{C_{0}}{x^{2}}\,\cos\!\bigl(2\sqrt{11}\,\ln x+\phi\bigr)\;+\;o(x^{-2})\,,\qquad x\to 0^{+}\,$,
with fitted constants
$C_{0}\approx 0.11325\,,\qquad\phi\approx-2.79\,.$
Refining the fit, we construct an explicit boundary approximation $y_{\mathrm{app}}$ such that
$y-y_{\mathrm{app}}=o(1)\,,\qquad x\to 0^{+}.$
Based on the Frobenius analysis below, we introduce the following approximation near $x=0$, retaining all terms up to order $o(1)$:
$y_{\mathrm{app}}(x)=\frac{1}{x^{2}}\Bigl[$ $\bigl(A_{0}+A_{1}x+A_{2}x^{2}\bigr)\cos\!\bigl(2\sqrt{11}\,\ln x\bigr)$ (10.2)
$+\bigl(B_{0}+B_{1}x+B_{2}x^{2}\bigr)\sin\!\bigl(2\sqrt{11}\,\ln x\bigr)\Bigr]\,,\qquad x\to 0^{+}\,.$
The fitted numerical coefficients are